40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 24 11:31:59 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling EX_fiber_MM.pdfEX_fiber_changes.png
    Reply  Tue Apr 24 18:54:15 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling 
       Reply  Wed Apr 25 19:09:37 2018, gautam, Configuration, ALS, New look EX Fiber coupling EX_fiber_MM.pdfIMG_6977.JPGIMG_6972.JPGIMG_6974.JPGIMG_6976.JPG
          Reply  Thu Apr 26 11:24:50 2018, gautam, Configuration, ALS, New look EX Fiber coupling - pol stability polStab.pdf
             Reply  Thu Apr 26 18:58:21 2018, Bruce, Configuration, ALS, New look EX Fiber coupling - pol stability 
             Reply  Fri Apr 27 01:36:02 2018, gautam, Configuration, ALS, IR ALS noise performance 2018_04_BeatMouth_POX.pdf2018_04_BeatMouth.pdfALSSpecgram.pdfALS_ASD.pdfALSstab.pdf
                Reply  Wed May 2 21:39:33 2018, gautam, Configuration, ALS, IR ALS for EY 
                   Reply  Fri May 4 21:17:57 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table 
                      Reply  Tue May 8 00:40:51 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table IMG_6986.JPGIMG_6987.JPG
                         Reply  Thu May 24 19:41:28 2018, gautam, Configuration, ALS, BeathMouth reinstalled on PSL table 
                      Reply  Thu May 24 13:06:17 2018, gautam, Configuration, ALS, DFD noises DFDcalib.pdfDFD_NB.pdf
                         Reply  Thu May 24 20:31:03 2018, gautam, Configuration, ALS, DFD noises TF_X_mag.pdfTF_X_phase.pdf
Message ID: 13889     Entry time: Thu May 24 19:41:28 2018     In reply to: 13824
Author: gautam 
Type: Configuration 
Category: ALS 
Subject: BeathMouth reinstalled on PSL table 

Summary:

  • DC light power incident on beat PD is ~400uW from the PSL and ~300uW from EX.
  • These numbers are consistent with measured mating sleeve and fiber coupler losses.
  • However, I measure an RF beatnote of 80mVpp (= -18dBm). This corresponds to a mode matching efficiency of ~15%, assuming InGaAs efficiency of 0.65A/W.

I find this hard to believe.

Details:

  • I took this opportunity to clean the fiber tips on the PSL table going into the BeatMouth.
  • PSL light power going into the BeatMouth is 2.6mW. Of which ~400uW reaches the Beat PD (measured using my new front panel monitor port).
  • Similarly, 1mW of EX light reaches the PSL table, of which ~300uW reaches the Beat PD.
  • The RF amplifier gain is 20dB, and RF transimpedance is 50 ohms.
  • Using the (electrical) 20dB coupled port on the front panel, I measured a beat signal with 8mVpp. So the actual beat note signal is 80mVpp.

Discussion:

As I see it, the possibilities are:

  1. My measurement technique/calculation is wrong.
  2. The beat PD is broken has optoelectronic different that is significantly different from specifications.
  3. The non-PM fiber lengths inside the beat box result in ~15% overlap between the PSL and EX beams. Morever, there is insignificant variation in the electrical beat amplitude as monitored on the control room analyzer. So there is negligible change in the polarization state inside the BeatMouth.

I guess #3 can be tested by varying the polarization content of one of the input beams through 90 degrees.

ELOG V3.1.3-